The difference between two numbers is 2, and the sum of their squares is 52. Find the numbers.
Answer by jorel555(1290)

Let the two numbers be n and n+2. Then:
n²+(n+2)²=52
n²+n²+4n-48=0
n²+2n-24=0
(n+6)(n-4)=0
n=4, -6
The numbers are 4 and 6. ☺☺☺☺